OpenStudy (anonymous):

Need help doing partial fraction decomposition step by step... 4/x(x-1) can someone please help

zepdrix (zepdrix):

We have two linear factors in our denominator, so our setup will look like this: \[\Large\rm \frac{4}{x(x-1)}=\frac{A}{x}+\frac{B}{x-1}\]Understand that part? :U

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

thanks for replying, my bad I left my computer unattended

hero (hero):

@LALFAN4E, do you have it from there?

hero (hero):

You still have to find the values of A and B so it's not done yet.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i solved that one

OpenStudy (anonymous):

i moved on to a harder question can you help me?

OpenStudy (anonymous):

1/[^{3}\]/\[^{3}-8\]

OpenStudy (anonymous):

disregard that it's 1/x^3-8

OpenStudy (anonymous):

Im at 1= (A+B)x^2 + (2A-2B+C)X + (4A-2C)

hero (hero):

You should learn how to post fractions properly. What you posted could be interpreted as \(\dfrac{1}{x^3} - 8\) rather than \(\dfrac{1}{x^3 - 8}\)

OpenStudy (anonymous):

yea my bad, how do you do that.. I'm pretty new

hero (hero):

To avoid the confusion, use parentheses to emphasize that x^3 - 8 is in the denominator: 1/(x^3 - 8)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ok I expanded the expression out... but as to solving for A B and C I'm confused... given that the total expression =1, what do I set A+B = ? then, 2A-2B+C and 4A-2C

OpenStudy (anonymous):

the examples in my book show expressions where x-3 = ...

OpenStudy (anonymous):

in other words... how do I form the system of equations to answer A B & C

zepdrix (zepdrix):

1/(x^3-8)? Hmm I don't remember how to take that one apart... I guess we can use the difference of cubes formula:\[\Large\rm =\frac{1}{(x-2)(x^2+2x+4)}\]

OpenStudy (anonymous):

got that...

OpenStudy (anonymous):

then expanded it for A B C

zepdrix (zepdrix):

\[\Large\rm =\frac{A}{x-2}+\frac{Bx+C}{x^2+2x+4}\]Like that?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

then cross multiplied

zepdrix (zepdrix):

\[\Large\rm 1=A(x^2+2x+4)+(Bx+C)(x-2)\]k

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

expand...

zepdrix (zepdrix):

Nahhh don't expand :\ I mean if you prefer to do it that way.. then fine..

OpenStudy (anonymous):

ok without expanding can you show me and explain?

zepdrix (zepdrix):

It's always nice to try and solve it without expanding if you're able. Just plug in key values for x. So in this case, x=2 would cancel out that entire B/C block of stuff, yes? So plugging in x=2 gives us,\[\Large\rm 1=A(2^2+2\cdot2+4)+0\]

OpenStudy (anonymous):

well [B(2)+C)] how is that gone?

zepdrix (zepdrix):

[B(2)+C)](2-2) Because it's multiplying the 2-2

OpenStudy (anonymous):

oh ... I see my bad. good

zepdrix (zepdrix):

So that gets us our A value nice and easy,\[\Large\rm A=\frac{1}{12}\]

zepdrix (zepdrix):

Let's plug it back into our setup:\[\Large\rm 1=\frac{1}{12}(x^2+2x+4)+(Bx+C)(x-2)\]

zepdrix (zepdrix):

Now our A is out of the way. Can you think of any number that might let us solve for B or C? :o We have another good one we can work with.

OpenStudy (anonymous):

gimme a sec, writing and trying to understand.

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

or am I completely off?

zepdrix (zepdrix):

Ah yes good :) x=0 will zero out the B unknown for us, allowing us to solve for C!

zepdrix (zepdrix):

x=0: \[\Large\rm 1=\frac{1}{12}(4)+(C)(-2)\]Mmm k what'd we get for C?

OpenStudy (anonymous):

got it!

OpenStudy (anonymous):

C= -1/3

zepdrix (zepdrix):

Ok good. Plug it back in,\[\Large\rm 1=\frac{1}{12}(x^2+2x+4)+\left(Bx-\frac{1}{3}\right)(x-2)\]Then to solve for B, plug in any other value. Preferably something that's easy to work with, like x=1.

OpenStudy (anonymous):

B=1/12 correct?

zepdrix (zepdrix):

Hmm I came up with -1/12. lemme check my math real quick.

OpenStudy (anonymous):

because B x -1 would be -B.... thus i get -B = -1/12

zepdrix (zepdrix):

So when you multiply everything out you should get,\[\Large\rm 1=\frac{7}{12}-B+\frac{1}{3}\]That look right? You distribute the -1 to the 1/3 also?

zepdrix (zepdrix):

the -1 to the -1/3 i mean*

OpenStudy (anonymous):

my bad.. i messed up... i had 9/12 -___- did the algebra inside wrong

zepdrix (zepdrix):

ah :3

zepdrix (zepdrix):

Cool I think we got it! Sheesh that one was tricky! :U

OpenStudy (anonymous):

it was man... really appreciate the help. this site is too damn good

OpenStudy (anonymous):

hey but wouldn't the answer come out to 10/12?

OpenStudy (anonymous):

oh nm.... dumb. I see my mistake again... -___-

OpenStudy (anonymous):

i'm stressed out, missed class lecture for the hardest section of my chapter test lol, but I think I got it from here on out

OpenStudy (anonymous):

Hey lastly man, this is the answer in the book...

OpenStudy (anonymous):

for B, why is there a (x+4) ?

zepdrix (zepdrix):

They just factored out the -1/12 from each term, if you distribute it back in, you can see what's going on, \[\Large\rm -\frac{1}{12}(x+4)=\color{orangered}{-\frac{1}{12}}x\color{royalblue}{-\frac{1}{3}}\] That matches what we determined for B and C, yes? \[\Large\rm \color{orangered}{B=-\frac{1}{12}}, \qquad \color{royalblue}{C=-\frac{1}{3}}\]